(b) Suppose currency to deposit rato is 040 and required reserve ratio is 0.10 and Govt. buys bonds from public which increase the Monetary Base by Rs. 10000. Find out how much money supply will be changed? Also find the money multiplier.
1
Expert's answer
2021-08-04T04:35:37-0400

Money supply "=10000\times (0.4+0.1)= 5000"

Money multiplier

"= \frac {1}{Reserve Ratio}"

"= \frac {1}{0.1}=10"